OpenStudy (anonymous):

when 15 is reduced by one-third of a number, the result is 12. What is the number?

OpenStudy (anonymous):

I got -9 is this correct?

OpenStudy (ash2326):

Check again there is a small mistake

OpenStudy (anonymous):

9?

OpenStudy (ash2326):

yeah, now you got it right ;)

OpenStudy (anonymous):

how is this 9 tho?

OpenStudy (anonymous):

If the answer were -9 then you would be subtracting a negative number, so instead of having 15-3=12 you would be doing 15-(-3)=18, which is not right. Do you see what I mean?

OpenStudy (anonymous):

15-(1/3)x=12, then u solve for x

OpenStudy (anonymous):

\[15-\frac{ 1 }{ 3 }n=12\]